If each pupil in a class sends a New Year card to every classmate , the total number of cards sent out will be 870 . Find number of pupils in the class ? Pls Help
Answer by checkley71(8403)

THE FORMULA FOR THIS PROBLEM IS:
N(N-1)=870
N^2-N-870=0
(N-30)(N+29)=0
N-30=0
N=30 ANSWER FOR THE NUMBER OF CLASSMATES.